An ohmic appliance is rated at 60 W when it is connected to a 120-V source.

A)If the power company cuts the voltage by 6 % to conserve energy, what is the current in the appliance?

Express your answer using two significant figures. I=____A

B)What is the power consumed by the appliance after the voltage drop? Express your answer using two significant figures. P=___W
Answer #1

Given

P = 60W

V = 120 volts


A) Let R is the resistance of appliance.

so, P = V^2/R

==> R = V^2/P

= 120^2/60

= 240 ohms

when voltage is cut by 6%, V' = 0.94*120

= 112.8 volts

I = V'/R

= 112.8/240

= 0.47 A

B) P = V*I

= 112.8*0.47

= 53 W